OpenStudy (anonymous):

How do I solve this using substitution? y=5x+2 y=y^2+x^2=40,000

hartnn (hartnn):

so where are you stuck? what have you tried?

OpenStudy (anonymous):

Pretty much the first step, when I combine them I'm not sure where to go afterwards

hartnn (hartnn):

there's no like terms to combine anyways, can you please check your 2nd equation? there are 2 = signs in it :P somethin's not correct...

OpenStudy (anonymous):

I was supposed to come up with the first equation myself so maybe thats the problem... and oh wow sorry I meant y^2+x^2=40000

OpenStudy (anonymous):

And when I said combine I mean substitute one of the equations into the other, unless I'm doing this whole lesson wrong :P

hartnn (hartnn):

no, you're not :) thats the correct approach! plug in y= 5x+2 in 2nd equation, what u get?

OpenStudy (anonymous):

So would that get rid of the y^2?

hartnn (hartnn):

it will get rid of y and you will have equation only in x !

hartnn (hartnn):

\(\Large (5x+2)^2 + x^2 = 40000\) can you solve that?

OpenStudy (anonymous):

I'm so confused, sorry! Are you saying there will be no y's in the new equation?

hartnn (hartnn):

ofcourse! because you replaced the 'y' with 5x+2

OpenStudy (anonymous):

Ohh I see! thats what I had written down but I thought it wasn't right :P I'll try to solve it

OpenStudy (anonymous):

So I'd simplify the (5x+2)^2 first, right? Which would be 25x^2+20x+4?

hartnn (hartnn):

correct now combine like terms

OpenStudy (anonymous):

45x^2+4?

hartnn (hartnn):

x^2 and x are not like terms!

hartnn (hartnn):

\(25x^20x+4+x^2 =40000\) are u sure its 40000? :P

hartnn (hartnn):

*** \(25x^2+20x+4+x^2 =40000\)

OpenStudy (anonymous):

Yes it is :P

OpenStudy (anonymous):

Do I subtract 4?

hartnn (hartnn):

no \((25+1)x^2 +20x + (4-40000) = 0\)

hartnn (hartnn):

thats how you combine like terms

hartnn (hartnn):

\(26x^2+20x-39996=0\) try to solve that

OpenStudy (anonymous):

+39996 on both sides?

hartnn (hartnn):

won't help thats a quadratic equation know any method to solve a quadratic equation?

OpenStudy (anonymous):

Ah I figured out what I was doing wrong, thank you for helping! :)